# ACE: A general-purpose non-Markovian open quantum systems simulation toolkit based on process tensors

### v1

@@ I Introduction
One of the most widely studied classes of open quantum systems is the spin-boson model, which describes (bio-)molecules [1, 2], resonant nanojunctions [19], as well as semiconductor nanostructures like quantum dots (QDs) [20]. The particular Gaussian character of the linear coupling to a bath of harmonic oscillators enables a treatment using path integrals [21], which has been the basis of, on the one hand, hierarchical equations of motion (HEOM) [22, 23] and, on the other hand, the iterative path integral scheme QUAPI [24, 25]. Both techniques have been further developed [26] and implemented in computer codes like HierarchicalEOM.jl [27] and PathSum [28], respectively.
@@
Only recently a new strategy to describe open quantum systems has emerged: the process tensor (PT) formalism [6, 29]. Here, environment influences are encapsulated and represented in efficient tensor network structures called process tensor matrix product operators (PT-MPOs) [30]. PT-MPOs can be constructed to depend only on the environment Hamiltonian and the system-environment interaction, and describe the impact of the environment irrespective of interventions performed on the system, such as unitary time evolution due to a time-dependent system Hamiltonian or measurements [6] [see Fig. 1(a)]. This has many advantages: First, the numerically challenging part, the PT-MPO calculation, has to be performed only once and the resulting PT-MPO can be reused many times, e.g., to optimize parameters or driving protocols for open quantum systems [31]. Second, the allowed interventions on the system include those needed to extract multi-time correlation functions. This is particularly useful for non-Markovian open quantum systems when the quantum regression theorem no longer holds [32]. Finally, a quantum system coupled to two or more environments can be simulated using two PT-MPOs calculated independently of each other. The result remains numerically exact [33, 30]. This fact can be used to investigate non-additive multi-environment effects [34, 35] [see Fig. 1(b)] as well as cooperative effects in multi-site quantum systems where each site is coupled to a local non-Markovian environment [36]. Due to the modularity and separation of concerns, PT-MPOs are promising for scalable schemes to simulate small to medium-sized quantum networks [37].
@@
The first algorithms to calculate PT-MPOs [29] started from a tensor network derived from path integrals [38], and were thus restricted to Gaussian environments like generalized spin-boson models. These are now implemented in the Python package OQuPy [39]. Subsequently, progress was made in two directions: First, new schemes provide orders-of-magnitude speed-up by employing a divide-and-conquer strategy [40] and periodic PT-MPOs [40, 41]. Second, algorithms for different [42] and more general [33] types of environments have become available. Specifically, in Ref. [33], we introduced the algorithm Automated Compression of Environments (ACE), which is applicable to any environment that can be described in terms of $N_{E}$ independent modes, such as phonon, photon, fermion, spin, and anharmonic environments. Moreover, the environment modes themselves may be subject to Markovian losses and they may be time-dependently driven. Changing the contraction order of the tensor network and employing an efficient ‘preselection’ scheme for the combination of PT-MPOs with large inner bonds yields a variant of ACE which is about one to two orders of magnitude faster [43].
@@ II.4 Outer bond reduction
Our strategy to deal with large outer bonds is to use the fact that in many situations there are several values of $\beta_{l}$ where $\mathcal{Q}^{\beta_{l}}_{d_{l},d_{l-1}}=0$ for all combinations of $d_{l}$ and $d_{l-1}$, or where several $\mathcal{Q}^{\beta_{l}}_{d_{l},d_{l-1}}=\mathcal{Q}^{\beta^{\prime}_{l}}_{d_{l},d_{l-1}}$ for different $\beta_{l}\neq\beta^{\prime}_{l}$. An example of the former is the case of the spin-boson model when the system coupling operator $\hat{A}$ in Eq. (9) is diagonal. Then, the PT-MPO matrices cannot directly induce transitions between system states and $\mathcal{Q}^{(\alpha_{l},\alpha^{\prime}_{l})}_{d_{l},d_{l-1}}\propto\delta_{\alpha_{l},\alpha^{\prime}}$, which reduces the number of values of the outer bond indices to at most $D^{2}$. The situation where PT-MPO matrices with different values of $\beta_{l}$ are identical has also been discussed for QUAPI simulations in Ref. [46]. Translating these dicussions to PT-MPO techiques corresponds to utilizing degeneracies of eigenvalues of the spin-boson system coupling operator $\hat{A}$. These degeneracies arise trivially when the environment is coupled only to one subsystem of a composite open quantum system, e.g., for quantum dots coupled to an optical microcavity as well as to a non-Markovian phonon bath [47, 48, 49]. Moreover, degeneracies appear when there are decoherence-free subspaces, e.g. in the case of the biexciton-exciton diamond-shaped four-level system in a quantum dot, where the two excitonic states with different spin selection rules couple identically to the local phonon bath [50, 51].
@@ IV.2 Usage of the ACE algorithm
for the Boson generator and equivalent with boson operators replaced by fermion operators for the Fermion generator. This contains several often-used models: The default value for Fermion_SysOp is |0><1|_2, which corresponds to the resonant-level model describing a particle number conserving hopping processes. The default value for Boson_SysOp is |1><1|_2, the projection onto the excited state, which describes a spin-boson model for a two-level system. The Jaynes-Cummings model is obtained by setting Boson_SysOp to |0><1|_2. Note that for the Boson generator automatically subtracts the polaron shift or reorganization energy [see discussion of Eq. (9)]. If this is not desired, it can be switched off by the command Boson_subtract_polaron_shift false.
@@ IV.3 Selection of methods
The Gaussian methods allow for memory truncation, i.e. neglecting the bath correlation function beyond $n_{\textrm{mem}}=t_{\textrm{mem}}/\Delta t$ time steps. The memory time can be set in the configuration file by the parameter t_mem. Note that one should set the memory cut-off $n_{\textrm{mem}}$ to a power of two for the divide-and-conquer as well as for the periodic PT-MPO method. Generally, the computation time does not monotonically decrease when the memory time is reduced. This is likely due to the fact that a sudden jump to zero in the effective bath correlation function results in spurious long-range temporal correlations that increase the inner bond dimensions of the PT-MPOs. Heurestically, we suggest starting with a value of $t_{\textrm{mem}}$ which is about a factor 4 longer than the time scale on which the bath correlation function is found to drop to zero by visual inspection, and then varying the memory time to find an optimum in computation time.
@@ IV.4 Convergence and fine tuning
Several convergence parameters exist that control the accuracy as well as the computation time of ACE and other PT-MPO techniques, such as the width of the time steps $\Delta t$ and MPO compression threshold $\epsilon$. As described above, these are set via parameters dt and threshold.
@@ IV.5 Trotter Errors
However, for simulations with more than one PT-MPO with mutually non-communiting interaction Hamiltonians, we suggest an alternative decomposition, where the order of applying the PT-MPO matrices (with respect to outer index multiplication) alternates. Here, the understanding of PT-MPO matrices as compressed environment propagators is useful [30]. For example, if there are two environment propagators $e^{\mathcal{L}_{E}^{(1)}\Delta t}$ and $e^{\mathcal{L}_{E}^{(2)}\Delta t}$ in addition to the system propagator $e^{\mathcal{L}_{S}\Delta t}$, the propagation over two time steps is given by
@@ IV.6 Example: Composite system of QD and microcavity
Note that the system initial state, Hamiltonians (pulses), Lindblandians, observables, and the system-environment coupling operator Boson_SysOp all act on the 6-dimensional composite system Hilbert space containing both QD and the truncated cavity mode. However, because the system-environment coupling operator is highly degenerate—which is automatically identified by the code—the PT-MPO calculation is only as difficult as that for an isolated two-level system.
The dynamics depicted in Fig. 4(b) can be understood in terms of adiabatic undressing [54]: The strong pulse leads to laser-dressing of states, such that (i) phonon-assisted transitions between dressed states are possible and lead to fast thermalization towards the lower dressed state, and (ii) the lower dressed states adiabatically evolves into the excited state as the pulse vanishes. A corresponding jump in the occupations of state $|e,0\rangle$ (excited states with zero photons in the cavity) is observed at the end of the pulse. The excitation is then transfered to the cavity via the QD-cavity coupling and then out-coupled from the cavity via cavity losses. Moreover, during the pulse, the dressing detunes the QD from the cavity frequency, which suppresses emission during the pulse and thereby reduces reexcitation. Because the laser is also off-resonant from the cavity, the phonon-assisted scheme combines several important features of single-photon generation: high single-photon purity, seperability of emitted photons from stray laser photons, and relatively high brightness. Because of these advantages, the phonon-assisted scheme is used in practical implementations of single-photon sources [55].
@@ IV.7 Example: Photon coincidences from superradiant QDs
Our final example illustrates the use of multiple PT-MPOs at the same time. Motivated by recent experiments, which demonstrated cooperative emission from indistinguishable QDs [56, 57, 58, 59], we consider two QDs each coupled to a local non-Markovian phonon bath and both QDs coherently coupled to the electromagnetic environment.
@@
$G^{(2)}(t,\tau)=\langle\sigma^{+}_{S}(t)\sigma^{+}(t+\tau)\sigma^{-}(t+\tau)\sigma^{-}(t)\rangle$
$=\textrm{Tr}\big{[}\big{(}\sigma^{+}(t+\tau)\sigma^{-}_{S}(t+\tau)\big{)}\big{(}\sigma^{-}_{S}(t)\rho(0)\sigma^{+}_{S}(t)\big{)}\big{]}$ (22)
are the unnormalized coincidences. From the last line of Eq. (22), it is clear that $G^{(2)}$ can be obtained by propagating the open quantum system over a time $t$, then applying operator $\sigma^{-}_{S}$ from the left and $\sigma^{+}_{S}$ from the right, propagating further over time $\tau$, and finally evaluating the observable $\sigma^{+}\sigma^{-}$. This is precisely how we evaluate $G^{(2)}$ using the ACE code with the configuration file in Fig. 5(b), where we chose $t=0$.

### v2

@@ I Introduction
One of the most widely studied classes of open quantum systems is the spin-boson model, which describes (bio-)molecules [1, 2], resonant nanojunctions [19], as well as semiconductor nanostructures like quantum dots (QDs) [20]. The particular Gaussian character of the linear coupling to a bath of harmonic oscillators enables a treatment using path integrals [21], which has been the basis of several practical methods. On the one hand, path integrals have been used to derive hierarchical equations of motion (HEOM) [22, 23], which are now a well established technique and implemented in several computer codes like QuTiP-BoFiN [24] and HierarchicalEOM.jl [25]. On the other hand, there are schemes to sum up the path integral approximately or exactly in ways that avoids the exponential scaling of the Feynman-Vernon expression with respect to the number of time steps $n$. Examples are the iterative path integral scheme QUAPI [26, 27], which reduces complexity by assuming a finite memory time. This method is only exponential in the number of time steps $n_{\textrm{mem}}$ within the memory time. The base of the exponential scaling can be reduced by the blip decomposition [28], which also provides a very quickly converging approximate summation scheme for nearly incoherent dynamics. The Small Matrix Path Integrals (SMatPI) [29] decomposition using a series of small matrices of the dimension of the squared system Hilbert space strongly reduces the memory requirements and is implemented in the code PathSum [30].
Recently, tensor network representations have been exploited for efficient open quantum system simulations. In the method TEMPO [31], the augmented density matrix of QUAPI is represented as a MPO. While tensor network representations permit direct contraction schemes like PC-TNPI [32], more commonly the key to their success lies in MPO compression [15], which has an established role in leading to a very efficient representation of one-dimensional structures. This general principle has also been transferred to other open quantum systems techniques like ML-MCTDH [33] and HEOM [34].
@@
At the core of our code is the process tensor (PT) formalism [6, 35], where environment influences are encapsulated and represented in efficient tensor network structures called process tensor matrix product operators (PT-MPOs) [36]. PT-MPOs can be constructed to depend only on the environment Hamiltonian and the system-environment interaction, and describe the impact of the environment irrespective of interventions performed on the system, such as unitary time evolution due to a time-dependent system Hamiltonian or measurements [6] [see Fig. 1(a)]. This has many advantages: First, the numerically challenging part, the PT-MPO calculation, has to be performed only once and the resulting PT-MPO can be reused many times, e.g., to optimize parameters or driving protocols for open quantum systems [37]. Second, the allowed interventions on the system include those needed to extract multi-time correlation functions. This is particularly useful for non-Markovian open quantum systems when the quantum regression theorem no longer holds [38]. Finally, a quantum system coupled to two or more environments can be simulated using two PT-MPOs calculated independently of each other. The result remains numerically exact [39, 36]. This fact can be used to investigate non-additive multi-environment effects [40, 41] [see Fig. 1(b)] as well as cooperative effects in multi-site quantum systems where each site is coupled to a local non-Markovian environment [42]. Due to the modularity and separation of concerns, PT-MPOs are promising for scalable schemes to simulate small to medium-sized quantum networks [43].
@@
The first algorithms to calculate PT-MPOs [35] started from a tensor network derived from path integrals [31], and were thus restricted to Gaussian environments like generalized spin-boson models. These are now implemented in the Python package OQuPy [44]. Subsequently, progress has been made in two directions: First, new schemes provide orders-of-magnitude speed-up by employing a divide-and-conquer strategy [45] and periodic PT-MPOs [45, 46]. Second, algorithms for different [47] and more general [39] types of environments have become available. Specifically, in Ref. [39], we introduced the algorithm Automated Compression of Environments (ACE), which is applicable to any environment that can be described in terms of $N_{E}$ independent modes, such as phonon, photon, fermion, spin, and anharmonic environments. Moreover, the environment modes themselves may be subject to Markovian losses and they may be time-dependently driven. Changing the contraction order of the tensor network and employing an efficient ‘preselection’ scheme for the combination of PT-MPOs with large inner bonds yields a variant of ACE which is about one to two orders of magnitude faster [48].
@@ II.4 Outer bond reduction
Our strategy to deal with large outer bonds is to use the fact that in many situations there are several values of $\beta_{l}$ where $\mathcal{Q}^{\beta_{l}}_{d_{l},d_{l-1}}=0$ for all combinations of $d_{l}$ and $d_{l-1}$, or where several $\mathcal{Q}^{\beta_{l}}_{d_{l},d_{l-1}}=\mathcal{Q}^{\beta^{\prime}_{l}}_{d_{l},d_{l-1}}$ for different $\beta_{l}\neq\beta^{\prime}_{l}$. An example of the former is the case of the spin-boson model when the system coupling operator $\hat{A}$ in Eq. (9) is diagonal. Then, the PT-MPO matrices cannot directly induce transitions between system states and $\mathcal{Q}^{(\alpha_{l},\alpha^{\prime}_{l})}_{d_{l},d_{l-1}}\propto\delta_{\alpha_{l},\alpha^{\prime}_{l}}$, which reduces the number of values of the outer bond indices to at most $D^{2}$. The situation where PT-MPO matrices with different values of $\beta_{l}$ are identical has also been discussed for QUAPI simulations in Ref. [51]. Translating these dicussions to PT-MPO techiques corresponds to utilizing degeneracies of eigenvalues of the spin-boson system coupling operator $\hat{A}$. These degeneracies arise trivially when the environment is coupled only to one subsystem of a composite open quantum system, e.g., for quantum dots coupled to an optical microcavity as well as to a non-Markovian phonon bath [52, 53, 54]. Moreover, degeneracies appear when there are decoherence-free subspaces, e.g. in the case of the biexciton-exciton diamond-shaped four-level system in a quantum dot, where the two excitonic states with different spin selection rules couple identically to the local phonon bath [55, 56].
@@ IV.2 Usage of the ACE algorithm
for the Boson generator and equivalent with boson operators replaced by fermion operators for the Fermion generator. This contains several often-used models: The default value for Fermion_SysOp is |0><1|_2, which corresponds to the resonant-level model describing a particle number conserving hopping processes. The default value for Boson_SysOp is |1><1|_2, the projection onto the excited state, which describes a spin-boson model for a two-level system. The Jaynes-Cummings model is obtained by setting Boson_SysOp to |0><1|_2. Note that the Boson generator automatically subtracts the polaron shift or reorganization energy [see discussion of Eq. (9)]. If this is not desired, it can be switched off by the command Boson_subtract_polaron_shift false.
@@ IV.3 Selection of methods
The Gaussian methods allow for memory truncation, i.e. neglecting the bath correlation function beyond $n_{\textrm{mem}}=t_{\textrm{mem}}/\Delta t$ time steps. The memory time can be set in the configuration file by the parameter t_mem. Note that one should set the memory cut-off $n_{\textrm{mem}}$ to a power of two for the divide-and-conquer as well as for the periodic PT-MPO method. Generally, the computation time does not monotonically decrease when the memory time is reduced. This is likely due to the fact that a sudden jump to zero in the effective bath correlation function results in spurious long-range temporal correlations that increase the inner bond dimensions of the PT-MPOs. Heuristically, we suggest starting with a value of $t_{\textrm{mem}}$ which is about a factor 4 longer than the time scale on which the bath correlation function is found to drop to zero by visual inspection, and then varying the memory time to find an optimum in computation time.
@@ IV.4 Convergence and fine tuning
Several convergence parameters exist that control the accuracy as well as the computation time of ACE and other PT-MPO techniques, such as the width of the time steps $\Delta t$ and MPO compression threshold $\epsilon$. As described above, these are set via parameters dt and threshold. Details of the convergence depend on the chosen method and are discussed in the respective method papers [35, 39, 45, 48]. Here, we summarize the main insights and general trends.
Figure 4: Convergence for a spin-boson model as shown in Fig. 3(f). Panels (a) and (b) depict compression error and computation time, respectively, as a function of the nominal compression threshold $\epsilon$. Simulations are performed using the tree-like contraction of Ref. [48] and with exponentially increasing thresholds from $\epsilon/10$ to $\epsilon$ (threshold_range_factor 10). The inset in panel (a) shows the Trotter error obtained by comparing the maximal difference in the observable between the best converged (with respect to $\epsilon$) simulation for a given $\Delta t$ and reference calculation with time step $0.025$ ps. The dashed gray line the indicates a $\mathcal{O}(\Delta t^{2})$ trend.
An example for convergence is shown in Fig. 4, which depicts compression error, Trotter error, and computation times for the spin-boson model example of Fig.  3(c) and (f). For these results, we modified the configuration file underlying Fig.  3(c): First, we selected the tree-like contraction scheme by adding a line use_combine_tree true, second, we adjusted the number of modes to a power of two Boson_N_modes 64 (as this is the optimal choice for the tree-like scheme) with fine tuning parameter threshold_range_factor 10 (see below). Finally, we proceeded to vary the parameters threshold and dt and recorded the overall run times on a conventional laptop computer with AMD Ryzen 7 5825U processor. Compression errors were obtained by comparing simulation results with larger threshold to results using the smallest threshold; dt was kept fixed in each such series of runs. Our definition of compression error is the maximal difference of occupations occurring at any time between $t=0$ and $t=t_{e}$. Trotter errors were obtained by comparing with the simulations with smallest threshold in each series.
@@
In the example in Fig. 4, we see that to achieve results converged to a Trotter error of about $10^{-3}$, a time step $\Delta t\lesssim 0.1$ ps is required. For $\Delta t=0.1$ ps, a threshold of the order $\epsilon\lesssim 10^{-5}$ yields a compression error comparable to the Trotter error. The corresponding calculations finish within a few seconds. Extrapolating the $\mathcal{O}(\Delta t^{2})$ trend, we estimate a Trotter error of $\sim 10^{-4}$ for $\Delta t\sim 0.025$ ps, for which a similar compression error is achievable within a few minutes.
@@ IV.5 Trotter Errors
However, for simulations with more than one PT-MPO with mutually non-communiting interaction Hamiltonians, we suggest an alternative decomposition, where the order of applying the PT-MPO matrices (with respect to outer index multiplication) alternates [^2] Trotter decompsitions with alternating order were also proposed, e.g., for the modular path integral method in Ref. [74]. Here, the understanding of PT-MPO matrices as compressed environment propagators is useful [36]. For example, if there are two environment propagators $e^{\mathcal{L}_{E}^{(1)}\Delta t}$ and $e^{\mathcal{L}_{E}^{(2)}\Delta t}$ in addition to the system propagator $e^{\mathcal{L}_{S}\Delta t}$, the propagation over two time steps is given by
@@ IV.6 Example: Composite system of QD and microcavity
Note that the system initial state, Hamiltonians (pulses), Lindblandians, observables, and the system-environment coupling operator Boson_SysOp all act on the 6-dimensional composite system Hilbert space containing both QD and the truncated cavity mode. However, because the system-environment coupling operator is highly degenerate—which is automatically identified by the code—the PT-MPO calculation is only as difficult as that for an isolated two-level system (see Sec. II.4 Outer bond reduction).
The dynamics depicted in Fig. 5(b) can be understood in terms of adiabatic undressing [60]: The strong pulse leads to laser-dressing of states, such that (i) phonon-assisted transitions between dressed states are possible and lead to fast thermalization towards the lower dressed state, and (ii) the lower dressed states adiabatically evolves into the excited state as the pulse vanishes. A corresponding jump in the occupations of state $|e,0\rangle$ (excited states with zero photons in the cavity) is observed at the end of the pulse. The excitation is then transferred to the cavity via the QD-cavity coupling and then out-coupled from the cavity via cavity losses. Moreover, during the pulse, the dressing detunes the QD from the cavity frequency, which suppresses emission during the pulse and thereby reduces reexcitation. Because the laser is also off-resonant from the cavity, the phonon-assisted scheme combines several important features of single-photon generation: high single-photon purity, seperability of emitted photons from stray laser photons, and relatively high brightness. Because of these advantages, the phonon-assisted scheme is used in practical implementations of single-photon sources [61].
@@ IV.7 Example: Photon coincidences from superradiant QDs
A further example illustrates the use of multiple PT-MPOs at the same time. Motivated by recent experiments, which demonstrated cooperative emission from indistinguishable QDs [62, 63, 64, 65], we consider two QDs each coupled to a local non-Markovian phonon bath and both QDs coherently coupled to the electromagnetic environment.
@@
$G^{(2)}(t,\tau)=\langle\sigma^{+}_{S}(t)\sigma^{+}_{S}(t+\tau)\sigma^{-}_{S}(t+\tau)\sigma^{-}_{S}(t)\rangle$
$=\textrm{Tr}\big{[}\big{(}\sigma^{+}_{S}(t+\tau)\sigma^{-}_{S}(t+\tau)\big{)}\big{(}\sigma^{-}_{S}(t)\rho(0)\sigma^{+}_{S}(t)\big{)}\big{]}$ (22)
are the unnormalized coincidences. From the last line of Eq. (22), it is clear that $G^{(2)}$ can be obtained by propagating the open quantum system over a time $t$, then applying operator $\sigma^{-}_{S}$ from the left and $\sigma^{+}_{S}$ from the right, propagating further over time $\tau$, and finally evaluating the observable $\sigma^{+}_{S}\sigma^{-}_{S}$. This is precisely how we evaluate $G^{(2)}$ using the ACE code with the configuration file in Fig. 6(b), where we chose $t=0$.
@@
IV.8 Example: Strongly structured environment
Figure 7: Simulation of the free decay of initial coherences in a two-level system that is coupled to an environment featuring the strongly structured spectral density of the FMO complex [67]. Panel (a) shows the configuration file and panel (b) the spectral density, while (c) and (d) depict the dynamics of the coherences for temperature $T=0$ and $T=300$ K, respectively. PT-MPO simulations with different thresholds are compared with the exact dynamics obtained from polaron transformation as well as with TEMPO simulations.
Finally, we consider an example that is particularly challenging for PT-MPO-based methods. Because the inner bonds of PT-MPOs account for the dynamics within the most relevant environment degrees of freedom [36], the bond dimension becomes large when the system is coupled to a large number of modes, which appear as a series of narrow peaks in a strongly structured spectral density. Such a situation arises, e.g., in the Fenna-Matthews-Olson (FMO) photosynthetic complex, where a spectral density consisting of 62 modes was estimated from experiments in Ref. [67]. Here, we use the same spectral density as Lorenzoni et al. in Ref. [68], where the 62 sharp $\delta$-like resonances are described by slightly broadened Lorentzians, and an Adolphs-Renger contribution is added to provide a low-energy background [see Fig. 7(b)].
We couple a spin-boson environment with this spectral density to a two-level system and observe the free decay of coherences. Specifically, the system is initialized in an equal superposition of ground and excited state, we adopt the trivial system Hamiltonian $H_{S}=0$, and we then extract the expectation value of the observable $\sigma_{x}$. The narrow peaks in the spectral density lead to long memory times, while at the same time the high frequencies of some of the modes necessitates resolving fast oscillations with small time steps. The most advantageous PT-MPO method is then the divide-and-conquer scheme of Ref. [45] without memory truncation. The corresponding configuration file is shown in Fig. 7(a).
The resulting dynamics of the free coherence decay is depicted in Fig. 7(c) and (d) for temperatures $T=0$ and $T=300$ K, respectively, and for different values of the threshold $\epsilon$ (requiring only the modification of the parameters Boson_temperature and threshold in the configuration file). The results are compared with the exact expression derived from polaron transformation [69] (which is available courtesy of our choice $H_{S}=0$)
[…]
At low temperature ($T=0$), the exact time evolution of the coherence shows an initial drop within the first $\sim 100$ fs, after which it fluctuates around a nearly constant value. The PT-MPO simulations match well with this behavior during the initial 100 fs but convergence out to larger times becomes very slow and challenging. Even for a relatively large threshold of $\epsilon=3\times 10^{-4}$ we find a PT-MPO with bond dimensions of $\sim 1000$, which entails a computation time of a few hours. Note that such a large bond dimension is not due to the particular method (nor its specific implementation) but must occur as an intrinsic feature of a PT-MPO: As PT-MPOs are entirely independent of the system Hamiltonian, they must be able to describe the environment response to all possible system drivings. Driving the system at a frequency that is resonant with any environmental mode leads to significant excitation of—and back-action from—that particular mode. Correspondingly, the inner bonds of the PT-MPO must accommodate the description of the large number of configurations of environment excitations that is reachable with arbitrary system driving.
By contrast, TEMPO [31] is a tensor network method that efficiently encodes temporal correlations within the state of the system for a given system Hamiltonian. This method is implemented in our framework with the binary TEMPO, which processes the same configuration files as ACE. The TEMPO results obtained with the configuration file shown in Fig. 7(a) match the exact expression in Eq. (23) perfectly. Due to the trivial system Hamiltonian $H_{S}=0$, temporal correlations in the states are simple, and the computation time is negligible. It should, however, be noted that for more general $H_{S}$ and environments with less structure, PT-MPO-based methods are typically significantly faster than TEMPO, see, e.g., examples in Ref. [35] and the supplementary material of Ref. [39]. Which method is more advantageous thus strongly depends on the details of the physical system at hand.
Finally, as shown in Fig. 7(d), for the free decay of coherences at physiological temperatures $T=300$ K, we find immediate agreement between the different methods. The computation times reflect the fact that the PT-MPOs have even larger bond dimensions compared with the respective simulation at zero temperature. However, because the coherences are strongly damped and decay to zero with a decay time of $\sim 60$ fs (obtained from an exponential fit), deviations to the exact result are also strongly suppressed. It should be noted that the overall error in general depends not only on the nominal threshold but also on the system Hamiltonian and on the chosen observable. For example, we observe that deviations from trace preservation eventually accumulate to several percent (not shown) even for the smallest threshold in PT-MPO simulations at $T=300$ K. Thus, TEMPO simulations also provides a useful reference in the high-temperature regime.